# What is a Schiff moment anyway?

### v1

@@ I.   Electron interaction with a nuclear charge distribution
$\begin{split}V_{E0}&=k\left(\int d^{3}R\,\rho_{n}\right)\left[\frac{1}{r}\right(ht]\\ V_{E1}&=-k\left(\int d^{3}R\,\rho_{n}\,X_{i}\right)\left[\frac{-x_{i}}{r^{3}}\right]\\ V_{E2}&=\frac{k}{2}\left(\int d^{3}R\,\rho_{n}\,X_{i}X_{j}\right)\left[-\frac{4\pi}{3}\delta_{ij}\,\delta(\vec{r})+\frac{3x_{i}x_{j}-r^{2}\delta_{ij}}{r^{5}}\right]\\ V_{E3}&=-\frac{k}{6}\left(\int d^{3}R\,\rho_{n}\,X_{i}X_{j}X_{\ell}\right)\left[-\frac{4\pi}{5}\delta_{\{ij}\,\nabla_{\ell\}}\delta(\vec{r})+\frac{3\delta_{\{ij}x_{\ell\}}r^{2}-15x_{i}x_{j}x_{\ell}}{r^{7}}\right].\end{split}$ (3)
@@ IV.   The $\mathcal{F}$-field
But electrons do not move classically inside atoms, so it is instructive to calculate the $\mathcal{F}$-field produced by an electron in a hydrogenic atom. Consider an electron in a superposition of nsns and $n^{\prime}p_{z}$ states, where $n,n^{\prime}$ denote hydrogenic principal quantum numbers: ${\left|{\psi}\right\rangle}=c_{s}{\left|{ns}\right\rangle}+c_{p}{\left|{n^{\prime}p_{z}}\right\rangle}$, with complex coefficients $c_{s}$ and $c_{p}$. The spatial wavefunctions of the atomic states are ${\left\langle{\vec{r}}|{ns}\right\rangle}=\mathfrak{R}_{ns}(r)Y_{00}(\theta\phi)$ and ${\left\langle{\vec{r}}|{n^{\prime}p_{z}}\right\rangle}=\mathfrak{R}_{n^{\prime}p}(r)Y_{10}(\theta\phi)$, where $\mathfrak{R}_{n\ell}(r)$ are hydrogenic radial functions and $Y_{\ell m}(\theta\phi)$ are spherical harmonics. Note that ${\left\langle{ns}\right|}\mathcal{F}_{z}{\left|{ns}\right\rangle}$ and ${\left\langle{n^{\prime}p}\right|}\mathcal{F}_{z}{\left|{n^{\prime}p}\right\rangle}$ both vanish, since $\mathcal{F}_{z}$ is an odd-parity operator sandwiched between definite-parity wavefunctions, which makes the expectation value integral equal zero. Therefore only the cross terms between $\left|{ns}\right\rangle$} and $\left|{n^{\prime}p}\right\rangle$} contribute, and so ${\left\langle{\psi}\right|}\mathcal{F}_{z}{\left|{\psi}\right\rangle}=2\,\mathrm{Re}(c_{p}^{*}c_{s}){\left\langle{n^{\prime}p}\right|}\mathcal{F}_{z}{\left|{ns}\right\rangle}$. Defining $\xi=\mathrm{Re}(c_{p}^{*}c_{s})$, the expectation value of the $z$-component of $\vec{\mathcal{F}}$ in the electron state $\left|{\psi}\right\rangle$} is

### v2

@@ I.   Electron interaction with a nuclear charge distribution
$\begin{split}V_{E0}&=k\left(\int d^{3}R\,\rho_{n}\right)\left[\frac{1}{r}\right]\\ V_{E1}&=-k\left(\int d^{3}R\,\rho_{n}\,X_{i}\right)\left[\frac{-x_{i}}{r^{3}}\right]\\ V_{E2}&=\frac{k}{2}\left(\int d^{3}R\,\rho_{n}\,X_{i}X_{j}\right)\left[-\frac{4\pi}{3}\delta_{ij}\,\delta(\vec{r})+\frac{3x_{i}x_{j}-r^{2}\delta_{ij}}{r^{5}}\right]\\ V_{E3}&=-\frac{k}{6}\left(\int d^{3}R\,\rho_{n}\,X_{i}X_{j}X_{\ell}\right)\left[-\frac{4\pi}{5}\delta_{\{ij}\,\nabla_{\ell\}}\delta(\vec{r})+\frac{3\delta_{\{ij}x_{\ell\}}r^{2}-15x_{i}x_{j}x_{\ell}}{r^{7}}\right].\end{split}$ (3)
@@ IV.   The $\mathcal{F}$-field
But electrons do not move classically inside atoms, so it is instructive to calculate the $\mathcal{F}$-field produced by an electron in a hydrogenic atom. Consider an electron in a superposition of $ns$ and $n^{\prime}p_{z}$ states, where $n,n^{\prime}$ denote hydrogenic principal quantum numbers: ${\left|{\psi}\right\rangle}=c_{s}{\left|{ns}\right\rangle}+c_{p}{\left|{n^{\prime}p_{z}}\right\rangle}$, with complex coefficients $c_{s}$ and $c_{p}$. The spatial wavefunctions of the atomic states are ${\left\langle{\vec{r}}|{ns}\right\rangle}=\mathfrak{R}_{ns}(r)Y_{00}(\theta\phi)$ and ${\left\langle{\vec{r}}|{n^{\prime}p_{z}}\right\rangle}=\mathfrak{R}_{n^{\prime}p}(r)Y_{10}(\theta\phi)$, where $\mathfrak{R}_{n\ell}(r)$ are hydrogenic radial functions and $Y_{\ell m}(\theta\phi)$ are spherical harmonics. Note that ${\left\langle{ns}\right|}\mathcal{F}_{z}{\left|{ns}\right\rangle}$ and ${\left\langle{n^{\prime}p}\right|}\mathcal{F}_{z}{\left|{n^{\prime}p}\right\rangle}$ both vanish, since $\mathcal{F}_{z}$ is an odd-parity operator sandwiched between definite-parity wavefunctions, which makes the expectation value integral equal zero. Therefore only the cross terms between $\left|{ns}\right\rangle$} and $\left|{n^{\prime}p}\right\rangle$} contribute, and so ${\left\langle{\psi}\right|}\mathcal{F}_{z}{\left|{\psi}\right\rangle}=2\,\mathrm{Re}(c_{p}^{*}c_{s}){\left\langle{n^{\prime}p}\right|}\mathcal{F}_{z}{\left|{ns}\right\rangle}$. Defining $\xi=\mathrm{Re}(c_{p}^{*}c_{s})$, the expectation value of the $z$-component of $\vec{\mathcal{F}}$ in the electron state $\left|{\psi}\right\rangle$} is